# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

### 09 Mar 2026 v1

@@ Abstract
We consider context-sensitive (binary) hypothesis testing for i.i.d. observations under a multiplicative weight function. We establish the logarithmic asymptotic, as the sample size grows, of the optimal total loss (sum of type-I and type-II losses) and express the corresponding error exponent through a weighted Chernoff information between the competing distributions. Our approach embeds weighted geometric mixtures into an exponential family and identifies the exponent as the maximizer of its log-normaliser. We also provide concentration bounds for a tilted weighted log-likelihood and derive explicit expressions for Gaussian and Poisson models, as well as further parametric examples.
Key words and phrases: hypothesis testing, weighted Chernoff information, exponential family, error exponent, context-sensitive loss.
2020 Mathematics Subject Classifications: 62F03, 60F10.
@@ 1 Introduction
where $\mathbb{P}$ and $\mathbb{Q}$ are probability measures on $\cal X$} dominated by a reference measure $\mu$. Without loss of generality one may take $\mu=\frac{1}{2}(\mathbb{P}+\mathbb{Q})$ and write $p=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mu}$ and $q=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mu}$. In the unweighted setting, the optimal sum of type-I and type-II error probabilities is characterized by $\mathrm{TV}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})$ and can be written as
$\int\limits_{{\mathcal{X}}^{n}}\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),\penalty 10000\ \penalty 10000\ p(x_{1}^{n})=\prod\limits_{i=1}^{n}p(x_{i}),\penalty 10000\ \penalty 10000\ q(x_{1}^{n})=\prod\limits_{i=1}^{n}q(x_{i}).$ (1.1)
@@
$D_{C}(\mathbb{P},\mathbb{Q})$ $:=-\ln\rho(p,q)=\max_{\alpha\in[0,1]}\bigl[-\ln\rho_{\alpha}(p,q)\bigr].$ (1.2)
@@
Chernoff also introduced an asymptotic efficiency notion for comparing two experimental designs $e=\frac{\ln\rho_{1}}{\ln\rho_{2}}$ such that $n$ observations on one test are equivalent (i.e., give asymptotically the same total loss as $n\to\infty$) to $en$ observations on another test; see [3].
The focus of this paper is on a context-sensitive (weighted) analogue of this quantity and its logarithmic asymptotics as $n\to\infty$, cf. [5, 6]. In the context-sensitive setting, a nonnegative weight function (WF) $\varphi(x_{1}^{n})$ is given that reweights the loss of a wrong decision depending on the realised sample. It indicates the importance (or irrelevance) of the sample $X_{1}^{n}$ for the given statistical problem.
Throughout the paper, we assume that the weight is compatible with the i.i.d. structure and factorises across observations (with a slight abuse of notation, we use the same symbol $\varphi$ for the one-step weight and its product extension).
@@
$\varphi({x}_{1}^{n})=\prod_{i=1}^{n}\varphi(x_{i}),\qquad\varphi\geq 0.$ (1.3)
Assumption 1.1 is crucial: it yields single-letter weighted affinities $\rho^{\mathrm{w}}_{\alpha}(p,q)=\int_{\cal X}\varphi(x)p(x)^{\alpha}q(x)^{1-\alpha}\mathrm{d}\mu(x)$ and leads to an additive error exponent in the logarithmic asymptotics.
Intuitively, in the Bayesian framework with priors $\pi_{0}=\pi_{1}=\frac{1}{2}$ [^1] Here $\pi_{0},\pi_{1}$ denote the prior probabilities of $H_{0},H_{1}$., the Bayesian risk equals one half of the total loss. In particular, for one observation the weighted (Bayesian) risk can be written as
$\frac{1}{2}\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x).$ (1.4)
Since $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$ for every $\alpha\in[0,1]$, we obtain for any $\alpha$
$\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\int\limits_{\cal X}\varphi(x)p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)=\rho^{\mathrm{w}}_{\alpha}(p,q),$
and hence $\int\varphi\min\{p,q\}\,\mathrm{d}\mu\leq\inf\limits_{\alpha\in[0,1]}\rho^{\mathrm{w}}_{\alpha}(p,q).$ Thus, one obtains the upper bound
$\int\limits_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\exp\{-D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\},$
@@
where $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q}):=\max\limits_{\alpha\in[0,1]}\bigl[-\ln\rho^{\mathrm{w}}_{\alpha}(p,q)\bigr]$ (see Definition 2.4 below). Under Assumption 1.1, the preceding bound factorises over $n$ i.i.d. observations and yields an upper bound of exponential order $\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\}$ (up to a constant factor). Theorem 3.1 shows that this exponential rate is sharp on the logarithmic scale.
Main result and contributions.
@@
Denote by $L_{n}^{*}$ the optimal total context-sensitive loss (sum of type-I and type-II losses, optimised over decision rules) for $n$ i.i.d. observations under Assumption 1.1. Our main theorem (Theorem 3.1) establishes an exact single-letter logarithmic asymptotic:
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\},\qquad n\to\infty,$ (1.5)
where the exponent is the weighted Chernoff information
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right].$ (1.6)
In particular, when $\varphi\equiv 1$, (1.6) reduces to the classical Chernoff information.
@@
We also extend the exponent characterisation to a finite family of simple hypotheses: the optimal $M$-ary exponent is governed by the minimum pairwise weighted Chernoff information (cf. [9] for the corresponding unweighted principle). A key ingredient of the proofs is an exponential-family representation of weighted geometric mixtures: we embed the path $\alpha\mapsto\varphi\,p^{\alpha}q^{1-\alpha}$ into a likelihood-ratio exponential family and identify the Chernoff exponent as a maximiser of its log-normaliser. This viewpoint is closely related to the unweighted treatment via likelihood-ratio exponential families and log-normalisers in [7], while the present work may be viewed as a weighted extension in which the base measure is tilted by the context weight. In addition, we derive non-asymptotic concentration bounds for tilted weighted log-likelihood ratios, and provide explicit closed-form expressions for $D_{C}^{\mathrm{w}}$ in several parametric models (see Section 4).
Related work.
Exponential error rates in hypothesis testing originate in Chernoff’s seminal work [3]; see also Hoeffding [10]. The context-sensitive (weighted) testing framework and the weighted information measures used here are developed in [5, 6]. For $\varphi\equiv 1$, Chernoff information admits an information-geometric interpretation via likelihood-ratio exponential families generated by geometric mixtures and their log-normalisers, see [7]. Our results may be viewed as a weighted extension in which the base measure is tilted by the context weight $\varphi$. For background on information geometry of exponential families, see [11]. Multiple hypothesis testing and the principle that the optimal exponent is governed by the minimum pairwise Chernoff information are discussed in [9].
Structure of the paper.
@@
Section 2 gives formal definitions of the weighted Bhattacharyya affinities and the weighted Chernoff information used throughout (Subsection 2.2; see Definitions 2.32.4). Section 3 states and proves the main asymptotic result (1.5) and develops the exponential-family/log-normaliser representation together with its information-geometric consequences. Subsection 3.4 studies the tilted weighted log-likelihood and derives non-asymptotic concentration bounds. Section 4 illustrates the general theory on Gaussian, Poisson, and exponential models, and includes an application to $M$-ary hypothesis testing (Subsection 4.4). Some auxiliary results are collected in the appendices.
@@ 2.1 Context-sensitive losses and weighted total variation
(and similarly for $E_{\varphi}(q)$). Throughout this section we assume that $E_{\varphi}(p),E_{\varphi}(q)\in(0,\infty)$. (Equivalently, $\rho^{\mathrm{w}}_{0}(p,q),\rho^{\mathrm{w}}_{1}(p,q)\in(0,\infty).$) Then $p^{*},q^{*}$ are probability densities and, under $\varphi(x_{1}^{n})=\prod\limits_{i=1}^{n}\varphi(x_{i})$, we have $\varphi(x_{1}^{n})p(x_{1}^{n})=E_{\varphi}(p)^{n}(p^{*})^{\otimes n}(x_{1}^{n})$ (and similarly for $q$).
@@
$\alpha_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{P}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,D(X_{1}^{n})\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,D(x_{1}^{n})\,p(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (2.1)
$\beta_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{Q}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,(1-D(X_{1}^{n}))\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,(1-D(x_{1}^{n}))\,q(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (2.2)
@@
$L_{n}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (2.3)
@@
Minimising pointwise over $D(x_{1}^{n})\in[0,1]$ yields $D_{n}^{*}(x_{1}^{n})=1$ when $p(x_{1}^{n})\leq q(x_{1}^{n})$ and $D_{n}^{*}(x_{1}^{n})=0$ when $p(x_{1}^{n})>q(x_{1}^{n})$, giving (2.3). ∎
@@
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}):=\frac{1}{2}\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,\big|p(x_{1}^{n})-q(x_{1}^{n})\big|\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (2.4)
@@
For $\varphi\equiv 1$ this reduces to the usual total variation distance. If $\varphi$ vanishes on a non-negligible set, ${\rm TV}_{\varphi}$ is, in general, a pseudo-distance; this is sufficient for our purposes since it characterises the weighted losses.
Using $\min\{a,b\}=\frac{1}{2}(a+b-|a-b|)$ in (2.3) and the definition of ${\rm TV}_{\varphi}$ yields
$L_{n}^{*}=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-{\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}).$ (2.5)

### 09 May 2026 v3

@@ Abstract
We study binary hypothesis testing for i.i.d. observations under a multiplicative context weight. For the optimal weighted total loss, defined as the sum of weighted type-I and type-II losses, we prove the logarithmic asymptotic
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\},\qquad n\to\infty,$
where $D_{C}^{\mathrm{w}}$ is the weighted Chernoff information. The single-letter form of the exponent relies on a structural assumption that the weight factorises across observations, $\varphi(x_{1}^{n})=\prod_{i=1}^{n}\varphi(x_{i})$; this restriction is essential for the single-letter representation and should be distinguished from the weaker qualitative description “multiplicative context weight”. The proof embeds the weighted geometric mixtures $\varphi p^{\alpha}q^{1-\alpha}$ into a likelihood-ratio exponential family and identifies the rate through its log-normaliser. We also derive concentration bounds for the tilted weighted log-likelihood, obtain closed forms for Gaussian, Poisson, and exponential models, and extend the exponent characterisation to finitely many hypotheses.
Keywords: hypothesis testing; weighted Chernoff information; weighted Bhattacharyya coefficient; exponential family; information geometry; context-sensitive loss.
2020 Mathematics Subject Classification: 62F03, 60F10.
Published version: Entropy 2026, 28(5), 536. DOI: 10.3390/e28050536.
@@ 1 Introduction
where $\mathbb{P}$ and $\mathbb{Q}$ are probability measures on $\cal X$} dominated by a reference measure $\mu$. Without loss of generality, one may take $\mu=\frac{1}{2}(\mathbb{P}+\mathbb{Q})$ and write $p=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mu}$ and $q=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mu}$. In the unweighted setting, the optimal sum of type-I and type-II error probabilities is characterized by $\mathrm{TV}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})$ and can be written as
$\int\limits_{{\mathcal{X}}^{n}}\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),\penalty 10000\ \penalty 10000\ p(x_{1}^{n})=\prod\limits_{i=1}^{n}p(x_{i}),\penalty 10000\ \penalty 10000\ q(x_{1}^{n})=\prod\limits_{i=1}^{n}q(x_{i}).$ (1)
@@
$D_{C}(\mathbb{P},\mathbb{Q})$ $:=-\ln\rho(p,q)=\max_{\alpha\in[0,1]}\bigl[-\ln\rho_{\alpha}(p,q)\bigr].$ (2)
@@
Chernoff also introduced an asymptotic efficiency notion for comparing two experimental designs $e=\frac{\ln\rho_{1}}{\ln\rho_{2}}$ such that $n$ observations on one test are equivalent (i.e., they give asymptotically the same total loss as $n\to\infty$) to $en$ observations on another test; see [1].
The paper studies a context-sensitive (weighted) analogue of this criterion and the logarithmic asymptotics of the optimal total loss as $n\to\infty$, in the framework of [3, 4]. In the weighted setting, a nonnegative weight function $\varphi(x_{1}^{n})$ reweights the loss of a wrong decision according to the realised sample. Thus, $\varphi$ acts as a context factor that changes the relevance of different observations for the statistical task.
Weights of this form arise naturally whenever observations are not equally informative for the inference task. Two canonical mechanisms produce such $\varphi$. In importance-type reweighting, samples drawn under a proposal density $g$ are used to perform inference with respect to a target $h$, and the Radon–Nikodym factor $\varphi(x)=h(x)/g(x)$ enters the loss as a strictly positive (non-indicator) tilt; this is the mechanism underlying the context-sensitive framework of [3, 4].
In applications, the informational value of an observation often depends on the underlying channel state. A canonical example, directly relevant to multiple hypothesis testing of transmission regimes, is a mobile communication channel modulated by a multi-zone coverage process (e.g., strong / weak / outage) along the receiver trajectory: samples acquired in outage carry little information about the regime and are weighted accordingly. Such reliability-weighted aggregation in multi-state channels was studied within multi-valued frameworks in [5, 6].
Under the standard assumption that the modulating state at time $i$ is determined by $X_{i}$ alone, the resulting weight is a strictly positive bounded function $\varphi(x)$ and extends multiplicatively to $X_{1}^{n}$. The weighted Chernoff information $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$ then quantifies the effective discrimination rate under channel-dependent reliability and reduces to the classical rate $D_{C}(\mathbb{P},\mathbb{Q})$ in the limit $\varphi\equiv 1$. Further parametric instances (Gaussian, Poisson, exponential) are worked out in Section 4.
Throughout we assume that the weight is compatible with the i.i.d. structure and factorises across observations; by abuse of notation, $\varphi$ denotes both the one-step weight and its product extension.
@@
$\varphi({x}_{1}^{n})=\prod_{i=1}^{n}\varphi(x_{i}),\qquad\varphi\geq 0.$ (3)
Assumption 1.1 is the key single-letter hypothesis. It yields the weighted affinities
$\rho^{\mathrm{w}}_{\alpha}(p,q)=\int_{\cal X}\varphi(x)p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),$
hence an additive logarithmic rate. For one observation and equal priors, the weighted Bayes risk equals
$\frac{1}{2}\int_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x).$
Since $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$ for every $\alpha\in[0,1]$,
$\int_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\rho^{\mathrm{w}}_{\alpha}(p,q),$
and therefore
$\int_{\cal X}\varphi(x)\min\{p(x),q(x)\}\,\mathrm{d}\mu(x)\leq\exp\{-D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\},$
@@
where $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q}):=\max_{\alpha\in[0,1]}\bigl[-\ln\rho^{\mathrm{w}}_{\alpha}(p,q)\bigr]$ (see Definition 2.4). Under Assumption 1.1, the same bound factorises over $n$ observations and yields the exponential scale $\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})\}$. Theorem 3.1 shows that this scale is exact on the logarithmic level.
1.1 Main Result and Contributions
@@
Let $L_{n}^{*}$ denote the optimal total context-sensitive loss (sum of weighted type-I and type-II losses, minimised over decision rules) for $n$ i.i.d. observations under Assumption 1.1. Our main theorem (Theorem 3.1) proves the single-letter logarithmic asymptotic
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\},\qquad n\to\infty,$ (4)
where the rate is the weighted Chernoff information
$D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right].$ (5)
For $\varphi\equiv 1$, (5) reduces to the classical Chernoff information.
@@
We also extend the exponent characterisation to a finite family of simple hypotheses: the optimal $M$-ary rate is the minimum pairwise weighted Chernoff information (cf. [7] in the unweighted case). A central technical device is an exponential-family representation of the weighted geometric mixtures $\alpha\mapsto\varphi\,p^{\alpha}q^{1-\alpha}$. This embeds the mixtures into a likelihood-ratio exponential family and identifies the exponent through the corresponding log-normaliser. We further derive concentration bounds for tilted weighted log-likelihood ratios and closed-form expressions for $D_{C}^{\mathrm{w}}$ in several parametric models; see Section 4.
1.2 Contributions
Items (N1)–(N4) below indicate new results; items (A1)–(A3) summarise definitions, geometric context, and tools adopted from the existing literature.
  • (N1)
    (New.) Theorem 3.1 establishes the logarithmic asymptotic (4) for the optimal weighted total loss under the factorised weight of Assumption 1.1, with rate given by the weighted Chernoff information (5).
  • (N2)
    (New.) The exponential-family representation of the weighted geometric mixtures $\alpha\mapsto\varphi\,p^{\alpha}q^{1-\alpha}$ (Section 3.2) and the resulting uniqueness of the optimal skewing parameter $\alpha^{\ast}$.
  • (N3)
    (New.) Concentration bounds for the tilted weighted log-likelihood and the finite-$n$ tail bound of Theorem 3.16 (Section 3.4).
  • (N4)
    (New.) Closed-form expressions for $D_{C}^{\mathrm{w}}$ in the Gaussian, Poisson, and exponential models (Section 4), and the $M$-ary extension showing that the optimal rate equals the minimum pairwise weighted Chernoff information.
  • (A1)
    (Adapted definitions.) The definitions of the weighted Bhattacharyya affinities and the weighted Chernoff information generalise the classical unweighted quantities of [1, 2] and follow the context-sensitive framework of [3, 4]; their asymptotic and information-geometric consequences developed below are new.
  • (A2)
    (Geometric context.) The information-geometric identities of Section 3.3 are derived in the spirit of the Chentsov–Amari–Nielsen framework [9, 8, 10, 11] but are stated and proved for the tilted log-normaliser $\hat{F}(\theta)=\ln\int\varphi(x)\,e^{\theta^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)$; the unweighted limit $\varphi\equiv 1$ recovers the classical statements of [12, 11].
  • (A3)
    (Standard tool.) The concentration argument uses the Azuma–Hoeffding/ McDiarmid inequality [13, 14]; the novelty lies in its application to the tilted weighted log-likelihood.
1.3 Related Work
The exponential theory of testing errors goes back to Chernoff [1] and Hoeffding [2]. The context-sensitive framework and the weighted information quantities used here were developed in [3, 4]. The information-geometric viewpoint on Chernoff information originates with Chentsov [9]; the dually flat structure of exponential and mixture families and the associated $\alpha$-divergences are developed in [8, 10], and the Chernoff point is characterised as the intersection of an exponential geodesic with the Kullback–Leibler bisector in [11]. For $\varphi\equiv 1$, the likelihood-ratio exponential family description is given in [12]; the present paper extends this picture to the tilted integrand $\varphi\,p^{\alpha}q^{1-\alpha}$. The minimum-pairwise principle for multiple testing is due to [7]. Weighting mechanisms for covariate-dependent relevance have also been studied outside the asymptotic error-exponent framework, e.g., adaptive-kernel conditional-independence testing [15].
1.4 Structure of the Paper
@@
Section 2 introduces the weighted Bhattacharyya affinities and the weighted Chernoff information. Section 3 proves the main asymptotic result (4) and develops the exponential-family and information-geometric identities. Subsection 3.4 studies the tilted weighted log-likelihood and derives finite-$n$ concentration bounds. Section 4 examines Gaussian, Poisson, and exponential models and includes the $M$-ary extension. Auxiliary computations are collected in the appendices.
@@ 2.1 Context-Sensitive Losses and Weighted Total Variation
(and similarly for $E_{\varphi}(q)$). Throughout this section, we assume that $E_{\varphi}(p),E_{\varphi}(q)\in(0,\infty)$. (Equivalently, $\rho^{\mathrm{w}}_{0}(p,q),\rho^{\mathrm{w}}_{1}(p,q)\in(0,\infty).$) Then $p^{*},q^{*}$ are probability densities and, under $\varphi(x_{1}^{n})=\prod\limits_{i=1}^{n}\varphi(x_{i})$, we have $\varphi(x_{1}^{n})p(x_{1}^{n})=E_{\varphi}(p)^{n}(p^{*})^{\otimes n}(x_{1}^{n})$ (and similarly for $q$).
@@
$\alpha_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{P}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,D(X_{1}^{n})\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,D(x_{1}^{n})\,p(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (6)
$\beta_{\varphi}(D)$ $:=\mathbb{E}_{\mathbb{Q}^{\otimes n}}\!\big[\varphi(X_{1}^{n})\,(1-D(X_{1}^{n}))\big]=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,(1-D(x_{1}^{n}))\,q(x_{1}^{n})\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (7)
@@
$L_{n}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\min\{p(x_{1}^{n}),q(x_{1}^{n})\}\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (8)
@@
Minimising pointwise over $D(x_{1}^{n})\in[0,1]$ yields $D_{n}^{*}(x_{1}^{n})=1$ when $p(x_{1}^{n})\leq q(x_{1}^{n})$ and $D_{n}^{*}(x_{1}^{n})=0$ when $p(x_{1}^{n})>q(x_{1}^{n})$, giving (8). ∎
@@
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}):=\frac{1}{2}\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\,\big|p(x_{1}^{n})-q(x_{1}^{n})\big|\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}).$ (9)
@@
For $\varphi\equiv 1$, this reduces to the usual total variation distance. If $\varphi$ vanishes on a non-negligible set, ${\rm TV}_{\varphi}$ is, in general, a pseudo-distance; this is sufficient for our purposes since it characterises the weighted losses.
Using $\min\{a,b\}=\frac{1}{2}(a+b-|a-b|)$ in (8) and the definition of ${\rm TV}_{\varphi}$ yields
$L_{n}^{*}=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-{\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n}).$ (10)